Luisa is 5 years younger than her sister Sigrid,8 years from now she will be four-fifths as old as her sister will be then. How old is each now?

Found 2 solutions by stanbon, mangopeeler07:
Answer by stanbon(75887)

Luisa is 5 years younger than her sister Sigrid,8 years from now she will be four-fifths as old as her sister will be then. How old is each now?
------------------------
Let Luisa be x yrs old now
The Sigrid is x+5 yrs old now
-----------------------------
Luisa is x+8 yrs old eight years from now.
Sigrid is (x+5)+8 eight years from now
EQUATION:
x+8 = (4/5)((x+5)+8)
5x + 40 = 4x + 52
x = 12 (Luisa's age now)
x+5 = 17 (Sigrid's age now)
==============================
Cheers,
Stan H.

Answer by mangopeeler07(462)

----"Luisa is 5 years younger than her sister Sigrid"----

Luisa=L
Sigrid=L+5

----"8 years from now she will be four-fifths as old as her sister will be then"----

L+8=4/5(L+5+8)

Combine like terms on the right side
L+8=4/5(L+13)

Distribute on the right side
L+8=4/5L+10.4

Subtract 8 from both sides
L=4/5L+2.4

Subtract 4/5L from both sides
1/5L=2.4

Multiply both sides by 5
L=12

Luisa=L
Sigrid=L+5

----"How old is each now?"----

Louisa=12
Sigrid=17